# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

### v7

@@ Abstract
Text-Based Person Search (TBPS) represents a critical engineering application of artificial intelligence (AI) in intelligent surveillance systems where automated person identification using natural language descriptions is essential. Current AI methods face significant limitations in achieving well-aligned cross-modal representations and capturing fine-grained correspondences essential for discriminating similar individuals. The fundamental challenge lies in developing intelligent systems that effectively process heterogeneous data sources where visual and textual features exist in different semantic spaces, creating substantial inter-modal gaps. To address these challenges, we propose Sew Calibration and Masked Modeling (SCMM), a novel AI framework that calibrates cross-modal representations through strategic techniques. SCMM introduces two innovative mechanisms: a sew calibration loss that dynamically aligns image-text features based on textual information quality using adaptive margin learning, and a masked caption modeling loss that establishes detailed cross-modal correspondences through transformer-based masked prediction tasks. The sew calibration component implements quality-aware adaptive constraints, enabling fusion strategies that accommodate varying information richness across modalities. While the masked modeling mechanism facilitates fine-grained feature learning by discovering detailed relationships between visual elements and textual descriptions through cross-modal attention. Our streamlined dual-encoder architecture maintains computational efficiency suitable for real-time applications while achieving superior performance through complementary alignment and correspondence learning strategies. Extensive experiments across three benchmark datasets demonstrate SCMM’s effectiveness, achieving superior Rank1 accuracy on three challenging benchmarks. The framework validates the practical applicability of quality-aware adaptive constraints and fine-grained correspondence modeling for real-world AI engineering applications.
keywords:
Text-Based Person Search, Cross-modal Learning, Person Re-identification, Metric Learning, Engineering Applications, Artificial Intelligence
journal: Engineering Applications of Artificial Intelligence
[…]
[label3]organization=Department of ECE, The University of British Columbia, city=Vancouver, postcode=V6T 1Z4, country=Canada
[…]
[label1]organization=Academy for Engineering &\&& Technology, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
[label4]organization=Department of Computer Science, University of Toronto, city=Toronto, postcode=M5S 1A1, country=Canada
[…]
[label2]organization=MEGVII Technology, city=Beijing, postcode=100096, country=China
[…]
[label7]organization=Faculty of Data Science, City University of Macau, city=Macau, postcode=999078, country=China
1 Introduction
Artificial intelligence (AI) applications in engineering have revolutionized modern computer vision and machine learning systems, driven by the urgent need to develop intelligent solutions that can process and integrate heterogeneous data sources for enhanced understanding and comprehensive analysis capabilities [1, 2]. Person re-identification (Re-ID) exemplifies a critical engineering application of AI, requiring sophisticated machine learning algorithms to process visual features across non-overlapping camera views for identifying target individuals in crowded scenes and enhancing security measures in smart city infrastructure [3]. However, traditional Re-ID approaches face significant operational limitations in real-world engineering scenarios when query images are unavailable, necessitating the development of alternative AI-driven solutions [4]. Text-Based Person Search (TBPS) addresses this critical engineering challenge by employing advanced artificial intelligence techniques to process textual descriptions alongside visual galleries, representing a significant advancement in practical AI applications for surveillance and security systems [5]. The implementation of multi-modal processing in TBPS enables the effective integration of rich semantic information from natural language descriptions with visual features from surveillance imagery [6], thereby providing intelligent solutions for Re-ID in scenarios where visual queries are unavailable or insufficient. Consequently, the paradigm shift from traditional visual-only approaches to sophisticated AI-powered cross-modal techniques addresses fundamental limitations in real-world deployment while opening new possibilities for intelligent surveillance and security applications [7].
@@ 1 Introduction
The effectiveness of AI-powered multi-modal systems depends critically on developing machine learning algorithms that can bridge the semantic gap between heterogeneous modalities while preserving discriminative information essential for accurate person identification [8]. Engineering challenges in TBPS stem from the fundamental differences in feature representations between textual and visual domains, where semantic concepts must be aligned across disparate embedding spaces using advanced AI techniques [9]. State-of-the-art AI architectures have emerged to address these engineering challenges, incorporating transformer-based attention mechanisms and deep learning models for cross-modal interactions to enhance intelligent processing capabilities [10]. Contemporary research in AI engineering applications emphasizes the development of sophisticated alignment strategies that can effectively combine complementary information sources while maintaining computational efficiency suitable for practical deployment in real-world systems [11]. The integration of these advanced AI techniques has proven essential for engineering applications requiring precise person identification with natural language queries, particularly in scenarios where traditional visual-only approaches prove insufficient for accurate target discrimination in intelligent surveillance systems [12].
@@
Compared to conventional Re-ID, the main objective of TBPS is to learn fine-grained cross-modal representations between visual and textual modalities [13]. As shown in Fig. 1, cross-modal representations have two key characteristics that contribute to effectively searching for a person from images: closely aligned cross-modal representations and fine-grained information correspondence. First, closely aligned cross-modal representations can effectively reduce the inter-modal gap, making it easier to locate specific persons using textual captions. Second, detailed correspondence is absolutely essential for TBPS, as fine-grained cross-modal information is often necessary to discriminate between similar persons [14].
In the TBPS domain, numerous approaches have been proposed from two distinct perspectives. First, some methods [15, 16] employ only dual encoders and align the two modal representations using symmetric loss functions. Although computationally efficient, these methods suffer from limited feature alignment capabilities. Alternative approaches [17] utilize multi-modal models with transformer-based [18] cross-attention mechanisms to improve cross-modal feature alignment and interaction. However, such methods often require high computational costs due to the exhaustive fusion of all possible image-text pairs during inference, hindering their efficiency in real-time applications. Second, to obtain robust fine-grained features, recent TBPS works have designed multi-level [19, 20, 21, 22], multi-granularity [23, 24] matching strategies, and specialized attention modules [8, 25, 26]. While these methods provide fine-grained features, they suffer from complex model architectures and expensive computations. Furthermore, the fine-grained features they produce are often limited and hinder significant performance improvements.
@@
    We propose an efficient and scalable framework for text-based person search, named SCMM, that calibrates cross-modal representations using a dual-encoder architecture and an auxiliary cross-modal decoder. The framework’s design effectively resolves the trade-off between the high computational cost of complex interaction models and the limited alignment capability of simpler methods, achieving superior performance with high inference speed suitable for real-world AI engineering applications.
  • 2.
    We introduce a novel sew calibration loss, which uniquely incorporates a quality-guided adaptive margin mechanism. By dynamically adjusting alignment constraints based on the information density of textual descriptions, our loss function more effectively bridges the inter-modal gap, pushing visual and textual feature distributions into a more compact and well-aligned shared embedding space.
  • 3.
    We propose an innovative MCM loss that establishes detailed cross-modal correspondences through masked prediction tasks and transformer-based cross-modal attention mechanisms. Our method facilitates fine-grained information integration by learning word-level relationships between visual elements and textual descriptions, enabling the framework to distinguish between similar individuals with subtle differences while maintaining computational efficiency through training-only decoder utilization.
  • 4.
    We conduct extensive experiments on three challenging benchmarks: CUHK-PEDES, ICFG-PEDES, and RSTPReID. SCMM demonstrates superiority by achieving new state-of-the-art Rank1 accuracies of $73.81\%$, $64.25\%$, and $57.35\%$ respectively. The establishment of new performance benchmarks not only validates the effectiveness of our fusion strategies for text-based person search, but also provides valuable baselines for future research in cross-modal AI applications.
@@
The structure of the remainder of this paper is outlined as follows. Sec. 2 reviews the related work on text-based person search, metric learning, and masked language modeling. Sec. 3 introduces the proposed SCMM method, including the sew calibration loss, masked caption modeling loss, and the overall framework. Sec. 4 presents the experimental setup, datasets, implementation details, and comprehensive experimental results, including comparison with state-of-the-art methods, ablation studies on different components, and qualitative analysis with visualization of attention maps and feature distributions. Sec. 5 concludes the paper with a summary of contributions and future research.
2 Related Work
2.1 Text-Based Person Search
@@
Text-based person search was first introduced by [8], which identifies person images in a gallery using only textual queries. Fundamentally, the task represents a critical application of multi-modal fusion, where the integration of heterogeneous data sources—visual and textual modalities—becomes essential for enhanced understanding and analysis capabilities. Early works utilized pre-processing methods to obtain external cues such as person segmentation [19] and human body landmarks [27]. The evolution towards sophisticated fusion architectures has become crucial, as multi-view fusion combining complementary information from multiple sources shows significant promise for enhancing prediction accuracy and robustness in cross-modal tasks [10].
@@
In recent years, end-to-end frameworks based on attention mechanisms [22, 24, 26] have become prevalent, addressing the fundamental challenge of semantic mismatch between features of different modalities. Cross-modal attention is critical for performing image-text interaction, where advanced fusion strategies must consider the unique attributes of cross-modal matching tasks [3]. Existing methods can be broadly classified into attention-explicit and attention-implicit approaches. Attention-explicit methods [2, 26] design specific attention modules according to multi-granularity and multi-level strategies, implementing early fusion techniques that combine features at the input level. For example, NAFS [14] conducts cross-modal alignments over full-scale features with a contextual non-local attention module, while CFine [25] utilizes cross-attention for multi-grained global feature learning, achieving impressive results through knowledge transfer from the CLIP [28] model.
In contrast, attention-implicit methods [22, 24] utilize transformer-based models with shared parameters to align cross-modal semantics implicitly, representing late fusion approaches that combine decisions from individual modality-specific models [10]. However, compared to performing in-depth cross-attention with task-driven approaches, existing methods face challenges in addressing information asymmetry problems and do not implement sufficient cross-interaction with generalized performance, highlighting the need for more sophisticated fusion techniques that enhance feature integration and establish deeper connections between global structure and local details [3].
2.2 Metric Learning
@@
Initially, metric learning used $L_{2}$ distance as the metric, with the goal is to minimize the $L_{2}$ distance between samples of the same class. Some $L_{2}$-based metric learning methods include Siamese Networks [29] and Triplet Networks [30]. With the development of deep learning, researchers started using softmax-based loss functions in metric learning to learn a more discriminative distance metric. Additionally, increasing the margin between classes is an intuitive approach to learn a better metric space [3]. L-Softmax [31] introduced the concept of margin on softmax function for the first time. The widely used CosFace [32] proposed large-margin cosine loss to learn highly discriminative deep features for face recognition. Circle loss [33] proposed a simple loss based on a unified loss for metric learning and classification. Recently, some works introduced adaptive margin into marginal loss [2]. They usually learn image quality implicitly and adjust the margin accordingly. In single-modal representation learning, they usually give large margins to high-quality samples for hard mining. Compared to them, we try to solve a cross-modal matching problem where samples from two modalities have different information volume. We give greater tolerance to less informative samples in TBPS.
However, existing metric learning approaches face significant limitations when applied to cross-modal scenarios. First, traditional methods primarily focus on single-modal optimization and fail to adequately address the heterogeneity gap between different modalities [34]. Second, fixed margin constraints cannot adaptively handle the varying information density across modalities, leading to suboptimal alignment performance. Finally, most current approaches lack fine-grained correspondence modeling capabilities essential for distinguishing between similar cross-modal samples [4].
2.3 Masked Language Modeling
MLM is a highly effective method for pre-training language models [35] by randomly selecting a certain percentage of words from the input sentence and then predicting the masked words based on the context of other words. Many cross-modal pre-training models have utilized MLM in their methods. For example, the work in [17] combines MLM with contrastive loss in the framework, which achieved impressive performance in their cross-modal tasks. The success of MLM in BERT [36] has proven its ability to adapt well to various downstream tasks, leading to generalized performance. In our fine-grained framework, the task-driven decoder utilizing masked caption modeling can facilitate generic cross-modal learning.
@@ 2.3 Masked Language Modeling
However, existing MLM-based cross-modal approaches face several limitations when applied to fine-grained person search tasks. First, traditional MLM methods primarily focus on intra-modal language understanding and lack explicit cross-modal interaction mechanisms, limiting their ability to establish detailed visual-textual correspondences [6]. Second, most existing approaches apply MLM uniformly without considering the varying information density and quality of textual descriptions, leading to suboptimal learning efficiency. Finally, current methods often require complex multi-stage training procedures or heavy computational overhead during inference, hindering their practical deployment in real-time applications [23].
3 Proposed Methods
@@
Formally, given a set of images with corresponding captions, denoted as $X=\{(I_{i},T_{i})\}^{N}_{i=1}$, where each image $I_{i}$ and its description text $T_{i}$ is associated with a person ID $y_{i}$. The text-based person search task aims to retrieve the most relevant Rank $k$ (e.g., $k=1,5,10$) person images efficiently from a large-scale gallery using textual queries. As an engineering application of AI, it requires developing machine learning algorithms that can bridge the semantic gap between visual and textual modalities while maintaining discriminative capabilities for accurate person identification in real-world scenarios. The fundamental engineering challenge lies in designing AI systems that can align heterogeneous feature representations across modalities, where visual features capture spatial and appearance characteristics while textual descriptions encode semantic attributes and contextual information. Effective strategies must address both global alignment between modalities and fine-grained correspondences that enable discrimination between similar individuals in practical applications.
@@
To address these engineering challenges in AI applications, we propose SCMM, a streamlined yet effective machine learning framework as illustrated in Fig. 2. Our approach implements a dual-encoder architecture using state-of-the-art AI models: ViT [37] and BERT [36] as backbone networks for visual and textual feature extraction respectively. The image encoder processes image patches from $I_{i}$ along with a vision classification (CLS) token, outputting an image feature sequence $\{v_{i}\}$ and a vision CLS token embedding $v_{i}^{c}$. Similarly, the text encoder generates a text feature sequence $\{t_{i}\}$ and a text CLS token embedding $t_{i}^{c}$ for caption $T_{i}$, following established AI engineering practices [14, 25]. Our framework integrates two complementary fusion mechanisms to calibrate cross-modal representations: the sew calibration loss with adaptive margin constraints for global alignment, and the masked caption modeling loss for establishing detailed cross-modal correspondences. The following subsections elaborate on each component. We first introduce the sew calibration loss that dynamically aligns cross-modal features based on textual information quality using machine learning techniques. Subsequently, we present the masked caption modeling mechanism that facilitates fine-grained information integration through transformer-based cross-modal attention and masked prediction tasks. The total loss function combines these complementary AI objectives to jointly optimize global cross-modal alignment and fine-grained correspondence learning, enabling the framework to achieve superior performance while maintaining computational efficiency suitable for real-world engineering applications through the streamlined dual-encoder architecture.
3.1 Sew Calibration Loss with Constraints

### v8

@@ Abstract
Text-Based Person Search (TBPS) aims to retrieve target person images from a large-scale gallery using natural language descriptions, posing fundamental challenges in cross-modal representation learning. Existing methods often struggle to bridge the semantic gap between heterogeneous modalities while capturing fine-grained correspondences essential for discriminating visually similar individuals. To address these challenges, we propose Sew Calibration and Masked Modeling (SCMM), a unified framework that calibrates cross-modal representations through complementary learning mechanisms. Notably, SCMM introduces two novel components: a sew calibration loss that dynamically aligns image-text features using quality-guided adaptive margins based on textual information density, and a masked caption modeling loss that establishes fine-grained cross-modal correspondences through transformer-based masked prediction. Additionally, the sew calibration mechanism implements bidirectional constraints to effectively compress same-identity features in the shared embedding space, while the masked modeling component leverages a cross-modal decoder to learn word-level visual-textual relationships, enabling discrimination of subtle attribute differences. Our dual-encoder architecture achieves an effective balance between representation capability and computational efficiency by employing a training-only decoder design. Extensive experiments on CUHK-PEDES, ICFG-PEDES, and RSTPReID benchmarks demonstrate that SCMM achieves state-of-the-art performance with Rank1 accuracies of 73.81%, 64.25%, and 57.35%, respectively. Comprehensive ablation studies validate the effectiveness of each proposed component.
I Introduction
Cross-modal representation learning has emerged as a fundamental research direction in deep learning, driven by the need to develop neural network architectures that can effectively bridge heterogeneous modalities for comprehensive understanding and reasoning [liu2025diverse, gong2024crossmodal, yang2025dynamic]. Person re-identification (Re-ID) exemplifies a challenging visual recognition task requiring deep neural networks to match individuals across non-overlapping camera views by learning discriminative visual representations [zhu2025improving, lu2023learning]. However, conventional Re-ID methods rely on visual queries, which limit their applicability when reference images are unavailable [ding2025dynamic]. Text-Based Person Search (TBPS) addresses this limitation by leveraging natural language descriptions as queries, necessitating the development of sophisticated cross-modal learning algorithms that can align visual and textual representations in a shared semantic space [isa, hao2025masked]. This task requires neural networks to learn robust multimodal embeddings that capture both global semantic alignment and fine-grained attribute correspondences [zhai2022lit], presenting unique challenges for representation learning in vision-language models [yang2025tienet].
@@ I Introduction
The effectiveness of vision-language models critically depends on developing learning algorithms that can bridge the semantic gap between heterogeneous modalities while preserving discriminative representations essential for accurate retrieval [gnarnn, peng2025deepfidelity, liu2024deeply]. The fundamental challenge in TBPS arises from the inherent differences between visual and textual feature spaces, where semantic concepts must be aligned across disparate embedding manifolds through effective neural network architectures [zhou2024blind, yu2025climbreid]. Recent advances in transformer-based architectures have shown promising results by incorporating cross-modal attention mechanisms that enable rich feature interactions between modalities [qian2025survey]. Contemporary research emphasizes developing efficient alignment strategies that can combine complementary information modalities while maintaining computational tractability for practical deployment [bai2023textbased]. These advances in deep learning have proven essential for cross-modal retrieval tasks requiring fine-grained discrimination, particularly in scenarios where subtle visual differences must be captured through textual descriptions [caibc].
@@
Compared to conventional Re-ID, the main objective of TBPS is to learn fine-grained cross-modal representations between visual and textual modalities [dssl]. As shown in Fig. 1, cross-modal representations have two key characteristics that contribute to effectively searching for a person from images: closely aligned cross-modal representations and fine-grained information correspondence. First, closely aligned cross-modal representations can effectively reduce the inter-modal gap, making it easier to locate specific persons using textual captions. Second, detailed correspondence is absolutely essential for TBPS, as fine-grained cross-modal information is often necessary to discriminate between similar persons [nafs].
In the TBPS domain, numerous approaches have been proposed from two distinct perspectives. First, some methods [dualpath, cmpmcmpc] employ only dual encoders and align the two modal representations using symmetric loss functions. Although computationally efficient, these methods suffer from limited feature alignment capabilities. Alternative approaches [alignbefore] utilize multi-modal models with transformer-based [attention] cross-attention mechanisms to improve cross-modal feature alignment and interaction. However, such methods often require high computational costs due to the exhaustive fusion of all possible image-text pairs during inference, hindering their efficiency in real-time applications. Second, to obtain robust fine-grained features, recent TBPS works have designed multi-level [vitaa, ssan, tipcb, safaNet], multi-granularity [mgel, ivt] matching strategies, and specialized attention modules [gnarnn, cfine, acsa]. While these methods provide fine-grained features, they suffer from complex model architectures and expensive computations. Furthermore, the fine-grained features they produce are often limited and hinder significant performance improvements.
@@
  • We propose SCMM, an efficient dual-encoder framework with an auxiliary cross-modal decoder that effectively balances high computational cost and alignment capability for scalable text-based person search.
  • We introduce a sew calibration loss with quality-guided adaptive margins for cross-modal metric learning, which dynamically adjusts alignment constraints based on textual information density to bridge the inter-modal gap.
  • We propose a masked caption modeling loss that establishes fine-grained cross-modal correspondences through masked prediction, enabling the learning of word-level visual-textual relationships for detailed discrimination.
  • We conduct comprehensive experiments on CUHK-PEDES, ICFG-PEDES, and RSTPReID, achieving state-of-the-art Rank1 accuracies of 73.81%, 64.25%, and 57.35%, respectively, validating SCMM’s effectiveness.
@@
The remainder of this paper is organized as follows. Sec. II reviews related work on text-based person search and metric learning. Sec. III details the proposed SCMM framework, including the sew calibration and masked caption modeling losses. Sec. IV presents the experimental setup, comprehensive results on three benchmarks, ablation studies, and qualitative visualizations. Finally, Sec. V concludes the paper with a summary of contributions and future research directions.
II Related Work
II-A Text-Based Person Search
@@
Text-based person search was first introduced by [gnarnn], which identifies person images in a gallery using only textual queries. Fundamentally, the task represents a critical application of multi-modal fusion, where the integration of heterogeneous data sources (visual and textual modalities) becomes essential for enhanced understanding and analysis capabilities. Early works utilized pre-processing methods to obtain external cues such as person segmentation [vitaa] and human body landmarks [verbalNet]. The evolution towards sophisticated fusion architectures has become crucial, as multi-view fusion combining complementary information from multiple sources shows significant promise for enhancing prediction accuracy and robustness in cross-modal tasks [qian2025survey].
@@
In recent years, end-to-end frameworks based on attention mechanisms [safaNet, ivt, acsa] have become prevalent, addressing the fundamental challenge of semantic mismatch between features of different modalities. Cross-modal attention is critical for performing image-text interaction, where advanced fusion strategies must consider the unique attributes of cross-modal matching tasks [xie2025multiscale]. Existing methods can be broadly classified into attention-explicit and attention-implicit approaches. Attention-explicit methods [gong2024crossmodal, acsa] design specific attention modules according to multi-granularity and multi-level strategies, implementing early fusion techniques that combine features at the input level. For example, NAFS [nafs] conducts cross-modal alignments over full-scale features with a contextual non-local attention module, while CFine [cfine] utilizes cross-attention for multi-grained global feature learning, achieving impressive results through knowledge transfer from the CLIP [clip] model. In contrast, attention-implicit methods [safaNet, ivt] utilize transformer-based models with shared parameters to align cross-modal semantics implicitly, representing late fusion approaches that combine decisions from individual modality-specific models [qian2025survey]. However, compared to performing in-depth cross-attention with task-driven approaches, existing methods face challenges in addressing information asymmetry problems and do not implement sufficient cross-interaction with generalized performance, highlighting the need for more sophisticated fusion techniques that enhance feature integration and establish deeper connections between global structure and local details [xie2025multiscale].
II-B Metric Learning
@@
Initially, metric learning used $L_{2}$ distance as the metric, with the goal of minimize the $L_{2}$ distance between samples of the same class. Some $L_{2}$-based metric learning methods include Siamese Networks [signature] and Triplet Networks [facenet]. With the development of deep learning, researchers started using softmax-based loss functions in metric learning to learn a more discriminative distance metric. Additionally, increasing the margin between classes is an intuitive approach to learn a better metric space [xie2025multiscale]. L-Softmax [lsoftmax] introduced the concept of margin on the softmax function for the first time. The widely used CosFace [cosface] proposed large-margin cosine loss to learn highly discriminative deep features for face recognition. Circle loss [circle] proposed a simple loss based on a unified loss for metric learning and classification. Recently, some works introduced adaptive margin into marginal loss [gong2024crossmodal]. They usually learn image quality implicitly and adjust the margin accordingly. In single-modal representation learning, they usually give large margins to high-quality samples for hard mining. Compared to them, we try to solve a cross-modal matching problem where samples from two modalities have different information volume. We give greater tolerance to less informative samples in TBPS. However, existing metric learning approaches face significant limitations when applied to cross-modal scenarios. First, traditional methods primarily focus on single-modal optimization and fail to adequately address the heterogeneity gap between different modalities [jiang2023crossmodal]. Second, fixed margin constraints cannot adaptively handle the varying information density across modalities, leading to suboptimal alignment performance. Finally, most current approaches lack fine-grained correspondence modeling capabilities essential for distinguishing between similar cross-modal samples [ding2025dynamic].
II-C Masked Language Modeling
@@
MLM is a highly effective method for pre-training language models [devlin2018bert] by randomly selecting a certain percentage of words from the input sentence and then predicting the masked words based on the context of other words. Many cross-modal pre-training models have utilized MLM in their methods. For example, the work in [alignbefore] combines MLM with contrastive loss in the framework, which achieved impressive performance in their cross-modal tasks. The success of MLM in BERT [lee2023lanobert] has proven its ability to adapt well to various downstream tasks, leading to generalized performance. In our fine-grained framework, the task-driven decoder utilizing masked caption modeling can facilitate generic cross-modal learning. However, existing MLM-based cross-modal approaches face several limitations when applied to fine-grained person search tasks. First, traditional MLM methods primarily focus on intra-modal language understanding and lack explicit cross-modal interaction mechanisms, limiting their ability to establish detailed visual-textual correspondences [zhai2022lit]. Second, most existing approaches apply MLM uniformly without considering the varying information density and quality of textual descriptions, leading to suboptimal learning efficiency. Finally, current methods often require complex multi-stage training procedures or heavy computational overhead during inference, hindering practical deployment in real-time applications [mgel].
III Proposed Methods
@@
Formally, given a set of images with corresponding captions, denoted as $X=\{(I_{i},T_{i})\}^{N}_{i=1}$, where each image $I_{i}$ and its description text $T_{i}$ is associated with a person ID $y_{i}$, the text-based person search task aims to retrieve the most relevant Rank $k$ (e.g., $k=1,5,10$) person images efficiently from a large-scale gallery using textual queries. To accomplish this, the task requires learning discriminative cross-modal representations that can bridge the semantic gap between visual and textual modalities while maintaining fine-grained correspondence for accurate retrieval. The fundamental challenge lies in designing neural network architectures that can align heterogeneous feature representations, where visual features capture spatial and appearance characteristics while textual descriptions encode semantic attributes and contextual information. Effective learning strategies must address both global alignment across modalities and local correspondences that enable discrimination between visually similar individuals.
@@
To address these challenges in cross-modal learning, we propose SCMM, a streamlined yet effective framework as illustrated in Fig. 2. Our approach employs a dual-encoder architecture using Vision Transformer (ViT) [vit] and BERT [lee2023lanobert] as backbone networks for visual and textual feature extraction, respectively. The image encoder processes image patches from $I_{i}$ along with a vision classification (CLS) token, outputting an image feature sequence $\{v_{i}\}$ and a vision CLS token embedding $v_{i}^{c}$. Similarly, the text encoder generates a text feature sequence $\{t_{i}\}$ and a text CLS token embedding $t_{i}^{c}$ for caption $T_{i}$, following established practices in vision-language learning [nafs, cfine]. Our framework integrates two complementary learning mechanisms to calibrate cross-modal representations: the sew calibration loss with adaptive margin constraints for global alignment, and the masked caption modeling loss for establishing fine-grained cross-modal correspondences. The following subsections elaborate on each component. We first introduce the sew calibration loss that dynamically aligns cross-modal features based on textual information quality. Subsequently, we present the masked caption modeling mechanism that facilitates fine-grained correspondence learning through transformer-based cross-modal attention and masked prediction. The total loss function combines these complementary objectives to jointly optimize global cross-modal alignment and fine-grained correspondence learning, enabling the framework to achieve superior performance while maintaining computational efficiency through the streamlined dual-encoder architecture.
III-A Sew Calibration Loss with Constraints